# A Reinforcement Learning Framework for Application-Specific TCP Congestion-Control

### v1

@@ Abstract
The Congestion Control (CC) module plays a critical role in the Transmission Control Protocol (TCP), ensuring the stability and efficiency of network data transmission. The CC approaches that are commonly used these days employ heuristics-based rules to adjust the sending rate. It is well known that due to their heuristics-based nature, these approaches are not only unable to adapt to changing network conditions but are also agnostic to the diverse requirements that different applications often have. Lately, we have seen several learning-based CC approaches that do not rely on fixed heuristics-based rules and thus can adapt to changing network conditions. Unfortunately, existing learning-based CC approaches have also not been designed to take application requirements into account. Prior heuristics-based as well as learning-based CC approaches focus on achieving a singular objective, which is often to maximize throughput, even though a lot of applications care more about latency, packet losses, jitter, and/or different combinations of various network metrics. Motivated by this, we propose a Deep Reinforcement Learning (DRL) based CC framework, namely ASC, which allows any application to specify any arbitrary objectives that the network traffic of that application should achieve and is able to swiftly adapt to the changes in the objectives of the applications as well as to the changes in the network conditions. Our ASC framework further employs a client-server architecture that serves two purposes: 1) it makes ASC highly scalable in terms of the arrival and departure of TCP connections, and 2) it makes ASC very lightweight for the nodes maintaining the TCP connections. We implemented and extensively evaluated ASC in a variety of settings. Our results show that it can not only achieve various objectives but also outperforms prior approaches even in the specific objectives that those approaches were designed to achieve.
I Introduction
I-A Background
@@ I-B Limitations of Prior Art
Application-agnostic: Different I-Apps have different needs. For example, file transfer needs high throughput [9], audio calls need low latency [10], cloud streamed games require low latency as well as small jitter [11], and so on. Unfortunately, as heuristics-based approaches rely on fixed rules, they cannot take into account such requirements of a diverse set of I-Apps simultaneously. Similarly, all existing learning based approaches fall short because they focus on a singular objective, such as maximizing throughput, and assume that all I-Apps have that same objective. It is also possible that an I-App’s objective changes over time, which, neither the heuristics-based nor the existing learning-based approaches handle.
@@
I-C Problem Statement
Our goal is to develop a congestion control framework that allows any I-App to specify any arbitrary objectives that the network traffic of that I-App should achieve, such as maximize throughput, minimize latency, minimize jitter, optimize any other network metric, or any combination of these objectives. The framework should further meet the following three requirements: (1) adapt to changes in the objectives of the I-App, (2) require minimal computational capability from the node running the I-App, and (3) be scalable in terms of swiftly handling new and terminating TCP connections.
I-D Proposed Approach
We present ASC, a reinforcement learning based congestion control framework that can achieve any application-specified objectives while meeting the three requirements listed above and addressing the two limitations of prior art. Logically, ASC performs two functions: learning and inference. Inference is computationally cheap and is performed on the node running the I-App(s). Learning, however, can be computationally heavy and, thus, ASC performs it on a dedicated server. Onwards, we will refer to the nodes performing the inference as CC-clients (i.e., congestion control clients) and the server performing the learning as CC-server.
@@ I-D Proposed Approach
Each CC-client periodically transmits some network statistics, called system state (details on the specifics of system state later), of each I-App running on it to the CC-server. Every time the CC-server receives the new-system state of any I-App from a CC-client, it updates the central-model as well as the sub-model of that I-App using reinforcement learning, and sends the updated inference-model back to the CC-client to be used for that I-App. Any CC-client receives inference-models only for the I-Apps executing on it. The CC-client then uses these inference-models to set the congestion window (cwnd)111 Congestion window (cwnd) is the tool that congestion control approaches use to adjust the amount of data that an I-App can send before it has to stop and wait for acknowledgement (ACK). Congestion control approaches mitigate congestion by appropriately limiting the amount of unACK’d data that is in-flight. Interested readers can learn more about cwnd here [12]. sizes for the TCP connections of the corresponding I-Apps. The use of these individualized inference-models ensures that the I-Apps achieve their specific objectives. The duration that the CC-client waits for while collecting system state before sending it to the CC-server and receiving the updated model from the CC-server warrants attention. A shorter updating interval results in increased resource consumption due to frequent communication with the server, while a longer interval hinders model’s performance by lacking latest information. In the evaluation section, we will thoroughly analyze the impact of updating interval on the network overhead as well as on ASC’s ability to achieve the specific objectives of the I-Apps.
@@
Adaptability: Our ASC framework can handle the scenarios where any given I-App updates/changes its objectives. When this happens, CC-server retrains the sub-model corresponding to that I-App, but not from scratch. ASC introduces a fast retraining mechanism (FRM), where instead of retraining from scratch, CC-server first determines whether any other I-App communicating through the same bottleneck link has a similar objective. If it identifies such an I-App, it uses the sub-model of that I-App as the starting model and adapts it for the I-App that just changed its objectives. This approach minimizes the need for extensive fine-tuning, thereby achieving significantly faster retraining. If the CC-server does not find another I-App with similar objective, it organically fine-tunes the sub-model of the I-App that just changed its objective. Note that the fine-tuning is still a fairly quick process because the sub-model is derived from a well-trained central-model.
Minimal Computational Requirements: In ASC, as inference is computationally very cheap, it is performed on the CC-client running the I-App(s). The computationally complex part is learning, which is performed on a dedicated CC-server. This deployment model makes ASC amenable even for low-powered devices.
@@
In household settings, the CC-server could be a typical computer within the house. It could also be a modern modem/router installed by the ISP in the house. The CC-server could even be provided as a cloud based service (we have not evaluated ASC with such a deployment of CC-server though). In enterprise settings, the CC-server could be deployed along with a DHCP server, a web server, an authoritative DNS server, an internet proxy, or any other such server. Similarly, in data centers, the CC-server could be deployed on one or more dedicated servers situated within that data center.
@@ I-E Key Contributions
We introduce FRM, which makes it practical for I-Apps to change their objectives and quickly achieve them.
@@
ASC is highly scalable in terms of swiftly adding new I-Apps and removing those that are not sending data anymore.
@@
We extensively evaluated ASC on multiple topologies and showed that it can not only achieve various objectives but also outperforms prior approaches even in the specific objectives that those prior approaches were designed to achieve.
@@ II Related Work
Heuristics-Based Congestion Control. Heuristic-based congestion control methods adjust the congestion window by utilizing predefined condition-action tables. In these methods, a TCP sender typically uses packet losses and round trip time (RTT) as indicators of the network conditions. TCP CUBIC [2] employs a cubic function of time since the most recent packet loss to adjust its cwnd. TCP-Illinois [13] aims to serve the needs of long-distance networks and relies on both packet losses and RTT to learn network conditions. FAST-TCP [14] aims to maintain a constant occupancy of the queue based on the RTT difference. Tan et. al. [15] found out that that FAST-TCP may lead to unfairness due to inaccurate RTT estimation and proposed an improved version. TCP Hybla caters to communication using terrestrial or satellite radio links. DCTCP [3] focuses on data center networks and employs explicit congestion notification to mitigate congestion. For a systematic investigation of the fairness and performance of some prominent TCP variants, interested readers can refer to [16, 17]. As mentioned earlier in Sec. I-B, unfortunately, due to the use of fixed rules, heuristics-based congestion control approaches are application agnostic and cannot handle specific requirements of the I-Apps.
Learning-Based Congestion Control. Unlike heuristics-based approaches, learning-based methods have the ability to learn policies from the environment rather than relying on predefined rules [18, 19]. Several works belonging to this category treat congestion control as an optimization problem and define utility functions [18, 19, 20, 21, 22]. Lately, the potential of reinforcement learning in congestion control has also been explored [23, 24, 5, 25, 26]. Li et. al. explored Q-Learning [5], while Tessler et al. investigated the policy gradient method [24] in data center networks. He et al. [27] introduced RMTC, capable of achieving real-time congestion control and multipath scheduling. In [27], multi-agent deep reinforcement learning and self-attention mechanisms were adopted for multipath congestion control. There are several other approaches that use similar learning-based techniques to address congestion, such as [28, 29, 30, 31, 26, 32]. Interested readers can find comparisons and analyses of the prominent learning-based congestion control methods in [23, 25, 33]. Unfortunately, all prior learning-based approaches suffer from the key limitations mentioned earlier in Sec. I-B. These include their inability to handle diverse optimization objectives from a diverse set of I-Apps, their inability to handle changes in the optimization objectives of the I-Apps, and their high computational demand from the clients executing these I-Apps.
@@ III ASC – Technical Details
We described the overall architecture of ASC in Sec. I-D. As the learning approach that ASC uses is the deep reinforcement learning (DRL), we first provide a quick primer of DRL. As we will see during the primer, the four components that we need to carefully design when using DRL include the System State, the Reward Function, the Action Space, and the Neural-Network Architecture. Therefore, after providing the primer, we will discuss how we design these four components of the DRL used in ASC.
@@ III-B System-State
Timestamp (1 value): Timestamps enable the model to learn any temporal patterns that show up repeatedly in the network traffic.
@@
RTT (2 values): Round-trip time (RTT) is the time it takes for a segment to go from the sender to the receiver node and its ACK to go from the receiver back to the sender. RTT is a crucial metric that quantifies the congestion on the path from the sender to the receiver, and is thus imperative for the DRL model to have. Similar to other metrics, a CC-client maintains the exponentially weighted value of RTT as well for any given I-App.
Throughput (2 values): Throughput and its exponentially weighted version represent the rate at which data is successfully delivered by the network from the sender to the receiver. They encompass information regarding both congestion and packet error rates, and thus provides the DRL model with a comprehensive overview of the path from the sender to the receiver.
Optimization Objective (1 value): This value identifies exactly which optimization objective any given I-App wants to achieve going forward. As soon as an I-App with existing TCP connection changes its optimization objective, or as soon as an I-App establishes a new TCP connection with a certain optimization objective, it informs the CC-server of that objective. If the CC-server has never seen that optimization objective previously, it assigns that objective a unique integer identifier and informs the CC-client running the I-App about that value. If the CC-server has seen that objective previously, it provides the CC-client with the integer identifier that it previously assigned to that objective. Subsequently, the CC-client uses this value as part of the system-state until either the TCP connection terminates or the I-App changes its objective.
@@ III-D Action Space
For any given I-App, at any given time step , the inference-model that the CC-client has for that I-App outputs a value of cwnd for use at time step . Another approach of determining the value of cwnd for the next time step would be to pre-define a step-size , use the model as a two class classification model, and add/subtract from the current cwnd value based on the output of the model. We went with the first approach because it makes the model more flexible and adaptive: the model is able to rapidly make large changes to cwnd or gradually make small changes to cwnd as needed in response to drastic or slow changes, respectively, in network conditions.
@@ III-F Fast Retraining Mechanism
Recall that an I-App has the liberty to change its optimization objective at any time. Typically, this would necessitate a complete retraining of its sub-model. We introduce fast retraining mechanism (FRM), aimed at substantially reducing the retraining latency. FRM works by leveraging the observations that in a given non-wide area network, if an I-App is running on one CC-client, it is quite likely that other instances of that I-App are running on other, or even the same, CC-client(s). Consequently, it is highly likely that some other instance(s) are already using the optimization objective that the current instance of the I-App just switched to. If it indeed is the case, the CC-server first identifies the corresponding existing sub-model for the I-App instance with the same objective. Next, it copies the 9664 weight matrix between Faltten and FC1 layers and the 6432 weight matrix between the FC1 and FC2 layers from the existing sub-model to the new sub-model. By employing this strategy of initialization, ASC enables new sub-model to effectively leverage past knowledge, drastically reducing adaptation time compared to complete retraining of the sub-model from scratch. Furthermore, we randomly initialize the 321 weight matrix between the FC2 and the output layers of the new sub-model instead of copying it from an existing sub-model. This enables the new sub-model to learn new information from the system state sent by its corresponding I-App instead of just completely copying everything from another sub-model related to a different instance of that I-App.
@@ III-G Scalability
An important property inherent in our design of ASC is scalability. This scalability stems from the simplicity with which ASC can handle new as well as terminating TCP connections. Whenever an I-App establishes a new TCP connection, the CC-server simply starts a new sub-model for this I-App and uses DRL to train over time. This training is quite fast because the central-model already exists and incorporates the latest information about the current status of the network. If the optimization objective of the new I-App matches with the optimization objective of another current I-App, the CC-server further employs FRM described above to speed up the training of the new sub-model for it to quickly start achieving that optimization objective. Similarly, if an I-App terminates its TCP connection, the CC-server simply discards its sub-model.

### v2

@@ Abstract
The Congestion Control (CC) module plays a critical role in the Transmission Control Protocol (TCP), ensuring the stability and efficiency of network data transmission. The CC approaches that are commonly used these days employ heuristics-based rules to adjust the sending rate. It is well known that due to their heuristics-based nature, these approaches are not only unable to adapt to changing network conditions but are also agnostic to the diverse requirements that different applications often have. Lately, we have seen several learning-based CC approaches that do not rely on fixed heuristics-based rules and thus can adapt to changing network conditions. Unfortunately, existing learning-based CC approaches have also not been designed to take application requirements into account. Prior heuristics-based as well as learning-based CC approaches focus on achieving a singular objective, which is often to maximize throughput, even though a lot of applications care more about latency, packet losses, jitter, and/or different combinations of various network metrics. Motivated by this, we propose a Deep Reinforcement Learning (DRL) based CC framework, namely ASC, which allows any application to specify any arbitrary objectives that the network traffic of that application should achieve and is able to swiftly adapt to the changes in the objectives of the applications as well as to the changes in the network conditions. Our ASC framework further employs a client-server architecture that serves two purposes: 1) it makes ASC highly scalable in terms of the arrival and departure of TCP connections, and 2) it makes ASC very lightweight for the nodes maintaining the TCP connections. We implemented and extensively evaluated ASC in a variety of settings. Our results show that it can not only achieve various objectives but also outperforms prior approaches even in the specific objectives that those approaches were designed to achieve. We will open-source the code and data used in the evaluation of ASC on the acceptance of this paper.
Index Terms:
TCP congestion control, deep reinforcement learning, user, specific congestion control, client-server updating mechanismI Introduction
@@ I-A Limitations of Prior Art
Application-agnostic: Different I-Apps have different needs. For example, file transfer needs high throughput [9], audio calls need low latency [10], cloud streamed games require low latency as well as small jitter [11], and so on. Unfortunately, as heuristics-based approaches rely on fixed rules, they cannot take into account such requirements of a diverse set of I-Apps simultaneously. Similarly, all existing learning based approaches fall short as they focus on a singular objective, such as maximizing throughput, and assume that all I-Apps have that same objective.
The objectives of many I-Apps of today change during different phases of their operation. For example, during the initial buffering phase, video streaming applications require high throughput to quickly fill the buffer to a sufficient level so that video playback starts right away. During the next playback phase, video streaming requires minimal jitter and packet loss to maintain smooth playback of the stream. During any re-buffering phase, video streaming again requires high throughput. Similarly, during web browsing, whenever a user clicks a link, low latency and high throughput are needed to minimize page load time. Afterwards, as the user interacts with the loaded page, throughput becomes less important and only low latency is needed to facilitate smooth interactions. Yet another example is video conferencing, where signaling/connection setup phase is sensitive to latency and jitter, video/audio streaming phase is additionally sensitive to packet loss, screen/content sharing phase is sensitive to throughput and latency, and remote screen control phase is sensitive to latency. A newer I-App, Cloud-streamed gaming, requires either low latency (to display immediate response when user provides inputs) or high throughput (to transfer video frames from cloud to the user’s device) and frequently switches between these two requirements as the user plays the game. There are numerous other such examples of I-Apps whose objectives change during different phases of their operations. Unfortunately, neither the heuristics-based nor the existing learning-based approaches can handle such changes in objectives.
@@
I-B Problem Statement
Our goal is to develop a congestion control framework that allows any I-App to specify any arbitrary objectives that the network traffic of that I-App should achieve, such as maximize throughput, minimize latency/jitter, optimize any other network metric, or any combination of these objectives. The framework should further meet the following three requirements: (1) adapt to changes in the objectives of the I-App, (2) require minimal computational capability from the node running the I-App, and (3) be scalable in terms of swiftly handling new and terminating TCP connections.
I-C Proposed Approach
We present ASC, a reinforcement learning congestion control framework that achieves any application-specified objectives while meeting the three requirements listed above and addressing the two limitations of prior art. Logically, ASC performs two functions: learning and inference. Inference is computationally cheap and is performed on the node running the I-App(s). Learning, however, can be computationally heavy and, thus, ASC performs it on a dedicated server. Onwards, we will refer to the nodes performing the inference as CC-clients (i.e., congestion control clients) and the server performing the learning as CC-server.
@@ I-C Proposed Approach
Each CC-client periodically transmits some network statistics, called system state (details on the specifics of system state later), of each I-App running on it to the CC-server. Every time the CC-server receives the new-system state of any I-App from a CC-client, it updates the central-model as well as the sub-model of that I-App using reinforcement learning, and sends the updated inference-model back to the CC-client to be used for that I-App. Any CC-client receives inference-models only for the I-Apps executing on it. The CC-client then uses these inference-models to set the congestion window (cwnd) sizes for the TCP connections of the corresponding I-Apps. The use of these individualized inference-models ensures that the I-Apps achieve their specific objectives. The duration that the CC-client waits for while collecting system state before sending it to the CC-server and receiving the updated model from the CC-server warrants attention. A shorter updating interval results in increased resource consumption due to frequent communication with the server, while a longer interval hinders model’s performance by lacking latest information. In the evaluation section, we will thoroughly analyze the impact of updating interval on the network overhead and on ASC’s ability to achieve the specific objectives of the I-Apps.
@@
Adaptability: Our ASC framework can handle the scenarios where any given I-App updates/changes its objectives. When this happens, CC-server retrains the sub-model corresponding to that I-App, but not from scratch. ASC introduces a fast retraining mechanism (FRM), where instead of retraining from scratch, CC-server first determines whether any other I-App communicating through the same bottleneck link has a similar objective. If it identifies such an I-App, it uses the sub-model of that I-App as the starting model and adapts it for the I-App that just changed its objectives. This approach minimizes the need for extensive fine-tuning, thereby achieving significantly faster retraining. If the CC-server does not find another I-App with similar objective, it organically fine-tunes the sub-model of the I-App that just changed its objective. Note that the fine-tuning is still a quick process because the sub-model is derived from a well-trained central-model.
@@
Minimal Computational Requirements: In ASC, as inference is computationally cheap, it is performed on the CC-client running the I-App(s). Learning is computationally complex, and thus performed on a dedicated CC-server. This deployment model makes ASC amenable even for low-powered devices. In household settings, the CC-server could be a typical computer in the house, or a modern modem/router installed by the ISP in the house. The CC-server could even be provided as a cloud based service (we have not evaluated ASC with such a deployment of CC-server though). In enterprise settings, the CC-server could be deployed along with a DHCP server, a web server, an authoritative DNS server, an internet proxy, or any other such server. Similarly, in data centers, the CC-server could be deployed on one or more dedicated servers situated within that data center.
@@ I-D Key Contributions
We introduce FRM, which makes it practical for I-Apps to change and quickly achieve their objectives.
@@
ASC is highly scalable in terms of swiftly adding new I-Apps and removing those that no longer send data.
@@
We extensively evaluated ASC on multiple topologies and showed that it not only achieves diverse objectives but also outperforms prior approaches, even in the specific objectives those approaches were designed for.
We will release the source code and data used in the evaluation and comparison of ASC on the acceptance of this paper.
@@ II Related Work
Heuristics-Based Congestion Control. Heuristic-based congestion control methods adjust the congestion window by utilizing predefined condition-action tables. In these methods, a TCP sender typically uses packet losses and round trip time (RTT) as indicators of the network conditions. TCP CUBIC [2] employs a cubic function of time since the most recent packet loss to adjust its cwnd. TCP-Illinois [12] aims to serve the needs of long-distance networks and relies on both packet losses and RTT to learn network conditions. FAST-TCP [13] aims to maintain a constant occupancy of the queue based on the RTT difference. DCTCP [3] focuses on data center networks and employs explicit congestion notification to mitigate congestion.
Learning-Based Congestion Control. Unlike heuristics-based approaches, learning-based methods have the ability to learn policies from the environment rather than relying on predefined rules [14, 15]. Several works belonging to this category treat congestion control as an optimization problem and define utility functions [14, 15, 16]. Lately, the potential of reinforcement learning in congestion control has also been explored [17, 5, 18]. Li et. al. explored Q-Learning [5], while Tessler et al. investigated the policy gradient method [19] in data center networks. He et al. [20] introduced RMTC, capable of achieving real-time congestion control and multipath scheduling. In [20], multi-agent deep reinforcement learning and self-attention mechanisms were adopted for multipath congestion control. There are several other approaches that use similar learning-based techniques to address congestion, such as [21, 18, 22]. Interested readers can find comparisons and analyses of the prominent learning-based congestion control methods in [17, 23, 24]. Unfortunately, all prior learning-based approaches suffer from the key limitations mentioned earlier in Sec. I-A. These include their inability to handle diverse optimization objectives from diverse I-Apps, their inability to handle changes in the optimization objectives of the I-Apps, and their high computational demand from clients executing these I-Apps.
Learning-Based Network Optimization. While not directly related, for the sake of completeness, we also mention some approaches that have used learning-based techniques to solve various network optimization problems. These problems include network data slicing [25, 26], load balancing [27], and resource management [28]. Luo et al. explored multipath data scheduling in [29], while Chinchali et al. [30] proposed a self-adaptive RL model for scheduling IoT traffic. Xie et al. [31] introduced a DRL-based method for adjusting the initial congestion window in 5G mobile networks. In [32], Nie et al. proposed TCP-RL, a DRL model that co-optimizes the initial window and congestion window. Li et al. [33] utilized GNNs to model network traffic, optimize flow routing, and perform topology management, resulting in reduced flow completion time. In [34], Almasan et al. investigated the combination of DRL with GNNs and demonstrated its potential through a routing optimization problem. For further details on GNNs in computer networks, refer to [35, 36].
@@ III ASC – Technical Details
We described the overall architecture of ASC in Sec. I-C. As the learning approach that ASC uses is the deep reinforcement learning (DRL), we first provide a quick primer of DRL. As we will see during the primer, the four components that we need to carefully design when using DRL include the System State, the Reward Function, the Action Space, and the Neural-Network Architecture. Therefore, after providing the primer, we will discuss how we design these four components of the DRL used in ASC.
@@ III-A DRL Primer
From the discussion above, it is evident that the information encapsulated in the system state, the reward, the action space, and the neural network, all play pivotal roles within the ASC framework. Next, we first describe these four components in detail. After that, we discuss a fast retraining mechanism that ASC uses and then conclude the section with a discussion on ASC’s scalability.
@@ III-B System-State
Timestamp (1 value): Timestamps allow the model to learn any temporal patterns that appear repeatedly in network traffic.
@@
RTT (2 values): Round-trip time (RTT) is the time it takes for a segment to go from the sender to the receiver node and its ACK to go from the receiver back to the sender. RTT is a crucial metric that quantifies the congestion on the path from the sender to the receiver, and is thus imperative for the DRL model to have. Similar to other metrics, a CC-client maintains the exponentially weighted value of RTT for any given I-App.
Throughput (2 values): Throughput and its exponentially weighted version represent the rate at which data is successfully delivered by the network from the sender to the receiver. They encompass information regarding both congestion and packet error rates and thus provide the DRL model with a comprehensive overview of the path from sender to receiver.
Optimization Objective (1 value): This value identifies exactly which optimization objective any given I-App wants to achieve going forward. As soon as an I-App with existing TCP connection changes its optimization objective, or as soon as an I-App establishes a new TCP connection with a certain optimization objective, it informs the CC-server of that objective.
@@ III-D Action Space
For any given I-App, at any given time step , the inference-model that the CC-client has for that I-App outputs a value of cwnd for use at time step . Another approach of determining the value of cwnd for the next time step would be to pre-define a step-size , use the model as a two class classification model, and add/subtract from the current cwnd value based on the output of the model. We went with the first approach because it makes the model more adaptive: the model is able to rapidly make large changes to cwnd or gradually make small changes to cwnd as needed in response to drastic or slow changes, respectively, in network conditions.
@@ III-F Fast Retraining Mechanism
Recall that an I-App has the liberty to change its optimization objective at any time. Typically, this would necessitate a complete retraining of its sub-model. We introduce fast retraining mechanism (FRM), aimed at substantially reducing the retraining latency. FRM works by leveraging the observations that in a given non-wide area network, if an I-App is running on one CC-client, it is quite likely that other instances of that I-App are running on other, or even the same, CC-client(s). Consequently, it is highly likely that some other instance(s) already uses the optimization objective that the current instance of the I-App just switched to. If it indeed is the case, the CC-server first identifies the corresponding existing sub-model for the I-App instance with the same objective. Next, it copies the 9664 weight matrix between Faltten and FC1 layers and the 6432 weight matrix between the FC1 and FC2 layers from the existing sub-model to the new sub-model. By employing this strategy of initialization, ASC enables new sub-model to effectively leverage past knowledge, drastically reducing adaptation time compared to complete retraining of the sub-model from scratch. Furthermore, we randomly initialize the 321 weight matrix between the FC2 and the output layers of the new sub-model instead of copying it from an existing sub-model. This enables the new sub-model to learn new information from the system state sent by its corresponding I-App instead of just copying everything from another sub-model related to a different instance of that I-App. In our current implementation, we do not rely on any form of pre-training on simulators. Instead, the central-model accumulates knowledge over time and each new sub-model benefits from it through parameter sharing. Bootstrapping from a pre-trained model may further enhance performance in early steps, and we leave that exploration to future work.
@@ III-G Scalability
An important property inherent in our design of ASC is scalability. This scalability stems from the simplicity with which ASC can handle new as well as terminating TCP connections. Whenever an I-App establishes a new TCP connection, the CC-server simply starts a new sub-model for this I-App and uses DRL to train over time. This training is quite fast because the central-model already exists and incorporates the latest information about the current status of the network. If the optimization objective of the new I-App matches with that of another current I-App, the CC-server further employs FRM described above to speed up the training of the new sub-model for it to quickly start achieving that optimization objective. Similarly, if an I-App terminates its TCP connection, the CC-server simply discards its sub-model.